Subject: APO 3.5m users committee minutes, March 2010
From: Suzanne Hawley
Submitted: Tue, 23 Mar 2010 23:17:06 -0700 (PDT)
Message number: 1174
(previous: 1173,
next: 1175
up: Index)
APO 3.5-m Users Committee Phonecon, 3/8/10
Attending: Suzanne Hawley, Mark Klaene, Bill Ketzeback,
John Bally, Michael Strauss, Jon Holtzman, Remy Indebetouw
Absent: Al Harper, Jon Fulbright, Scott Anderson
**********************************
User feedback, comments from institutional representatives:
New Mexico State (Jon Holtzman) - nothing to report (too cloudy).
Colorado (John Bally) - nothing to report
Princeton (Michael Strauss) - nothing to report
Virginia (Remy Indebetouw) - nothing to report, users happy.
Washington (Suzanne Hawley) - a beta version of TUI that allows
Agile on-chip guiding is about ready to be tested by UW users.
Others who are interested in testing should contact Suzanne.
Also, re: cloudy weather, OPEN and DD time can be requested for
critical programs that get weathered out (e.g. for graduate student
thesis data). Quite a bit of the OPEN/DD time has gone to such
programs in the past few months.
Chicago (Al Harper) - no report (absent)
Johns Hopkins (Jon Fulbright) - no report (absent)
*********************************
Discussion of telescope/instruments report - Klaene
There was some discussion of snow removal. We use
several methods including the snow program, roof heaters,
and manual shoveling. Which is used depends on conditions, powder
verses ice and wet snow. The snow program works well for powder, while
wet snow requires sun and manual shoveling. Ice on the lightning
arrestors has hurt us twice this month. Danger of this ice falling
into the open shutter is our biggest concern and there is no way to
manually remove the ice since the arrestors are quite tall and located
at the corners of the roof. Usually the ice clears with a few hours
of sunshine and/or ambient temps in the high 20s. The roof heaters
have limited value and can cause other problems.
Added 5 new beds (singles) in a few dorm rooms to accomodate
extra housing demands due to reduced availability of housing at
Sunspot. Sunspot/NSO housing is likely going away completely by
summer because new development folks working on ATST are
spending a few years at Sunspot before going to Hawaii.
The telescope should not be slewed with the covers closed. It is still
okay to do cals with covers closed while the telescope is tracking.
Bally asked about 2010 weather statistics. From SDSS they reported loss
time in January due to weather at 55%, which is quite a bit higher
than usual for the winter months.
******************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 2/3/10
through 3/3/10
0) Overview
Weather, specifically snow, has been the main issue again this period
and winter is not quite over. Recent storms have been more moderate
and caused significantly less problems although we continue to spend
over 2 man days a week dealing with weather related issues. This
now includes 3 roof leaks. Loss time due to snow on the roof
increased somewhat because of icing of the roof mounted lightning
arrestors. The arrestors are impossible to reach and clear and thus
we must have some warming to clear these. The concern with ice on
the arrestors is that it can easily fall into the open shutter and
onto the optics.
Remodeling of the dormitories was accomplished to add 5 more beds.
1) Telescope
Behavior has been quite good with only rare minor issues with
pointing and tracking. A couple of servo errors were seen during
slews and tracked to a software bug in the controller in handling
magnasensor crossings. Slewing with the mirror covers closed also
causes rare servo errors, we have not seen any problems while
tracking with the mirror covers closed. This resulted in a minor
procedure change.
2) Instruments
Echelle and Agile ran with no problems.
DIS blue warmed unexpectedly, and SPICam was running warming than
normal by 10-20 deg K. Vacuum was checked on both cameras and found
to be OK. Both cryotigers were recharged to a higher pressure based
on discussions with the Polycold technicians. The new pressure is
higher than their manual suggested for these ambient temperatures but
appears to have fixed the problem.
TSpec ICC power supply failed, taking out numerous boards in the
chassis. An SDSS spare chassis was assembled and UVa folks reinstalled
the software returning TSpec to normal operation. Fortunately the
PCIA Interface was okay. The spare is at CU. New chassis has been
procured and we plan to replace the power supply in the slitviewer ICC.
The DIS slitviewer camera shutter continues to have occasional
cold weather problems. We are monitoring this for increasing
frequency but so far it has not re-occurred. Observing Specialist are
running the shutter prior to start of observing when possible.
The new NICFPS fanout board testing is progressing at CU. Initial
testing with the spare Leach controller yielded unexpected results.
They have moved to the Rockwell controller and obtained more
consistent data. Once testing with the Rockwell controller is
complete they will revisit the issue with the spare Leach controller.
Installation is still possible for March or April. There was still no
progress reported on the NICFPS etalon at CU.
3) CIF projects
The spare drive box machining and fabrication is in work. A
preliminary vibration study was also conducted and is being analyzed.
We plan to perform additional testing this month. On chip guiding
software development for Agile is progressing and initial testing is
planned for later this month.
**********************************
o Q2 schedule and class trips - Hawley
The Q2 schedule was just released yesterday (3/7/10). There are no
class trips planned for Q2 according to the proposals that were
scheduled. If anyone is planning a large trip, they should let
us know immediately. Also, we need to be notified about trips that
are planned for Q3 (July-Sept) as soon as possible, since we will
be coordinating with summer shutdowns and SDSS work on the mountain.
o TUI versions - Klaene
There was a question about when the version of TUI being used
by an observer needs to be updated. The plan is that a new
field will be added that will come up when TUI is started so
that the version is available for the user and the Observing Specialist
to see. If it is out of date, this field will appear in some
bright color (possibly red). In general, users should try to
update their version of TUI to the newest version within a few
months of its release. The version history available at:
http://www.apo.nmsu.edu/35m_operations/TUI/index.html
gives the updates. Note that the current (latest) version is
TUI 1.7.0. Changes in the second digit (currently "7")
indicate a fairly major update that should be installed quickly,
while changes in the third digit (currently "0") are minor updates
often for a particular instrument. The new python libraries needed
for this update from 1.6.2 to 1.7.0 can be somewhat problematic to
install on linux machines.
o Possible visiting instrument: RITMOS - Hawley
RITMOS is a multi-object spectrograph built as a prototype instrument
for a 24" telescope and is presently available to go on a bigger telescope.
It would need some new front end optics. The FOV on the 3.5m would likely
be quite small, about 2'x1', similar to Agile and the GFP. We would like
to know if users would be interested. There is a published SPIE paper
(Meyer et al 2004 SPIE 5492, 200) that describes the instrument which has
been sent to the users committee members. Massimo Robberto at STScI is
the person who is leading this effort, because of his interest in
building a multi-object spectrograph, perhaps for the 3.5m, that uses this
technology.
o Note added in proof - Hawley
After the meeting finished, a new problem was discovered with the
SPIcam cryotiger. The instrument will be offline until May.
Updates will be sent to apo35-general.
*****************************
ACTION ITEMS:
[open from previous months]: NONE
[new actions from this meeting]:
Users committee members should ask their users if RITMOS is of interest
to them (please share SPIE paper).
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 5 April,
at 8:30 AM Pacific Time. The agenda and other materials will be sent
to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1174 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO